A report was made about a connection that couldn't be made on the S-string of the connection.
The connections failed to connect to the backend of the network and entire failed to connect to the V'R'S to the connection server.
Update: 05:17 am before closing the task. The connections seem stable, and we will monitor the progress of the situation.
Closed by ScoobyMedia - FREEDAR!
We had reports of addresses going elsewhere within the networks.
The CST teams have intervened to re-correct the problem.
DNS webmail assessed [OK]
Due to the ongoing security, we are doing security patches on our networks.
Your servers will auto reboot and will send you an auto message once the systems have rebooted.
Reason for closing: Done!
Hello, (Plesky) ( Backup ) services
The backup process will be backing up on 11-06-2022 GMT Date: Time 04:45 AM.
ALL the websites and MYSQL will back up in case of MYSQL fails!! ( We always keep MYSQL ) backup for the customer's safety and if we need to roll back an update or patch. We can roll back at a safe level.
SSL failed on cdncloud.scoobynetworks.co.uk we are working on a fix.
Please don't open tickets we are fully aware of this problem. We will provide more updates shortly!
we are gonna be improving the network so services will be offline for a while
we will update when we have completed the updates
00:16 start of the intervention
SNG-Host Network Teams
The manager v2 has a known bug in the Sonic Panel API functions when you click on your product/services.
We are working on a fix.
The radio access control panel is fixed.
Open Internal Ticket: Increase PHP
Intervention is done > Global PHP 7.4.29 settings were successfully applied to domains
HP max_execution_time 300 Minimum 600 is recommended
Good Evening Customers
Tonight our network team will be doing an emergency Security Patch.
This will be started at midnight UK time and will last roughly and hr we will update when this has been done and completed.
Pesky running some updates on Wordpress don't be alarmed if your website refreshes. This normal
Thank you SNG-HOST
we was made aware of a issue with the plesk module the customers use to login into their plesk control panel this was going to the port number causing the load time to increase.
We can no confirm that this issue has been rectified and a fix has been implemented.
Any issue dont hesitate to contact us
We currently looking further into mail issues concerning our mail services.
Please bare with us whilst we look further into this
Scheduled Maintenance tonight at 12 midnight
there will be some emergency security patches
What Should I Expect?
Although the scheduled maintenance window is 1 hour. We expect any downtime to be 60 minutes
During the Updates all websites and servers will be offline
If you require any further information the Please contact our Customer care Department
Scooby Networks Group CST
[Scheduled] Studio Database/CDN Upgrades (26th September)
Why are we doing this?
Upgrading the database improves stability and scalability of the platform. The newer database platform offers 8x the performance of our current infrastructure.
What should I expect?
Although the scheduled maintenance window is 4 hours. We expect any downtime to last up to 60 minutes.
During the upgrade, your station may fall silent while services are restarted. SNG-Host provided players may also cease to load. Live broadcasts should not be affected.
If you require any additional information or help, please contact our support team via email email@example.com
You can track the process here: https://scoobynetworks.uk/accounts/serverstatus.php
We thank you for your patience during this time.
Radio Develops Teams!
Good Afternoon Customers
We are aware of a issue when customers go to show live DJ widget on the djpanel.
We have escalated this issue to our CST teams and they are working to get this error fixed
When we have more information we shall update this task
Kind Regards Scooby Networks Group Customer Care
UPDATE: We have no fixed the issue with the widget on the DJ panel this is now working
SSL ISSUES / POP3 connection error can't connect to host error = stream_socket_client unable to connect to SSL mail.**
Unknown error ERRNO=0
- Stream_socket_client is able now to connect to SSL mail**
- We are checking the backup manager to see if the network is processing the updates.
- A process is running we are checking see if task finishes correctly.
- 88% backup holding and transferring to the CLOUD no further in-put's are needed at this time. We will check and monitors the current backups
- Current operation Server Status (OK) Current operation Server (Backup) (OK) Overall progress 100%
Furthermore, we are checking the timeout process for the current status of the backups. We unable to add any more information at this time.
Current operation The backup task is stopping. The process hanged was unable to stop, we ended the process we are checking further.
The Backup Manager is 100% we are checking since we have made a fix.
The backup task 100% is now fixed.
We made a change to the firewall to allow the sync of the clock for Microsoft only UDP connection.
These changes didn't have no effect on the whole-network.
The reason for closing the open case: No further actions need to be taken.
UPDATE: WE HAVE HAD TO RE-OPEN THIS DUE TO FURTHER ISSUES OUT OF SNG HOST'S CONTROL WE WILL UPDATE WHEN WE HAVE FURTHER INFORMATION.
We made some changes to the firewall we checking see if this fixed the current issues on 09/09/21 If you have any issues, please don't open support request we are aware of the current problem. We will provide updates very soon.
We are aware of an issue with POP3 Error: last request failed: Message 27 expunged.
We are looking into this and will update as soon as we can.
Update: We are checking further into the problem. We are unable to provide any support on this problem at the present moment.
On the 9/9/21 We are updating you on the current open request for the POP3 error: This still on-going we have fixed some of the issues we will remain looking further into the current issue provide more updates. The mail - servers are working status is normal.
Thanks CST - Customer's Care Department!
We are aware of an issue with customers getting a 502 error bad gateway.
We have passed this onto our NOC and are awaiting more information.
Once we have this we will update this notice.
NOC: We confirm the issue was caused by metadata that is blocked. We are still looking further into the current problem. The 502 error caused by word-press plugin. We made some changes to the following firewall this shouldn't be triggered by the mod-security either.
UPDATE: this issue has now been fixed any issue feel free to contact Customer Care
SNG Customer Solutions Teams
We are checking the mail server upon the connections at this time.
The connection limit in place. We are checking if causing anymore issues with the webmail the error message : internal error to server should now be fixed.
UPDATE: THIS ISSUE HAS NOW BEEN RESOLVED
Good Evening We are investigating a issue with Plesk at this moment in time we dont have any further information, We are awaiting more information Directly from Plesk themselves.
We will update you more when we have more information.
We have found fix this moment we still checking make sure the fix is working with out causing any issues.
Scooby Networks Group CST (Customer Solutions Teams)
We are fully aware of 502 bad gateway.
Our CST is looking further into the issue. We will update once we have more details.
We have updated the Ruleset. We check further into it see if the error 502 bad gateway fixed.
We are still seeing 502 bad gateway errors we are checking further into the cause of this problem we hope have fix soon.
this is Regarding "Freedar" and does not affect Scooby Networks Customers.
tonight our CST team have began the process of moving Freedar Servers from windows to linux. We shadowed the Image and uploaded these to a new server.
over the next few nights we will be moving servers.
There will be some affect to the radars this will affect the DESKTOP and MOBILE APP. This will be from 12am GMT
Should you face any issues please dont hesitate to contact us on firstname.lastname@example.org
Freedar Network Teams
IP ADDRESS: timestamp error this was reported.
We checked the services all normal without any errors.
The back-end process, we checked the hardware firewall found fault this was passed over to the NOC teams.
We updated the IP address ***** (This was blocked due to the Administrator) Due to security reason the information was removed due to data protection.
IP address routing was found to be the issue we re-profiled the IP address to the firewall services no longer disconnection or issues with timestamp error.
The intervention was closed and no further action is taken.
We are doing some backups on the networks/services - connection might be slow connection to the websites. We sorry about this we had no options to doing new backups when this settings was missing in the new update. We will update this status message once things return back to normal!
Kind Regards, CST
The Process of the backups on the entire network was done closed 10:26pm
The following Internal Ticket was opened due to change on the PHP Values.
Task in process now!
We checking see if able to increase the ratelimit or the limit in place in the API requests to make the control-panels more response to customer input?
Checked we can make improvements to the code. We have added code to inprove the API requests we checking for results.
The -rlimit increased all okay closed!
Dear Customer's We are aware of an issue with Plesk not loading
We are investigating this issue and will update you all when done
We have lost access to Plesk control panel this DOES NOT affect the customers, this is our side, we are working on this and will update you all when this has been fixed.
UPDATE:- Dear Customers: We are glad to inform you that the issue has now been resolved we have applied a hot fix
Should you have any issues feel free to open a support ticket with customer care?
Applying security measures to WordPress Domain.
The Wordpressy bots in action to-night working around the Word Pressy :)
Updating the WordPress plugins: Done
Updating the WordPress themes: Done
Reason for closing: Done!
We are in the process mitigating the data back to our Cloud Plesk.
Please bare with us whiles these tasks in process our PLESKY BOTS working around the clock.
We was notified this morning about a issue with the Website and Plesk login.
Our Network Engineers have rectified this issue and everything is back online.
Should you still face any issues please contact Customer Care by opening a Support Request Via your Customer Manager.
Scooby Networks Customer Solutions
We are just checking the RSS feeds.
Update: We have update our RSS feeds.
The DMARC - Email, DNS records showing the problem in the config we are trying resolve.
A ticket internal we are waiting on reviews.
Our records show this problem now fixed.
The current bot's are in force updating security on Wordpress sites this due to ensure you remain safe.
Heads up: Your website's outdated - and at risk of being hacked
Our current bots in force to ensure things run smoothly
We can confirm the IP profile for the Anti-DOS systems.
The systems are update this was rolled out to the entire network.
If you're having issues connecting to your services since these changes, please contact Customer Care.'
Great news! We would like to inform you all we are working on Smart Anti-DOS program again.
We don't have any ETA on when the process will start or change.
We are making some changes in the coming weeks, so we are processing backup's.
Once these backups done, we will close down the internal job.
The current back-up process done: We are moving files to our safe-place.
We are in the process of adding new firewall rules into our network
We have Added IP Profile's to all our IP addresses as we seen large amount attacks by passing the Anti-Vac systems.
Good morning. We are currently updating our Web host services & Mod-Security with latest patches. This protected our websites the connection to our control panels or the online websites may run slow or degrade in the connection speed. This normal whiles this work is carried out. We will update you once the updates* finished if you require support in the meanwhile, please contact our customer care by phone or support tickets. Thank you Administrator. (Happy morning)
UPDATE: This has been completed no further issues Thank you CST
Our engineer found fault with the radio services today on the 7th Feb 2021.
The case was un-known we have restarted the services that failed to start.
We closed this intervention down since this no longer fault with the service.
We checking further into the current issue with webmail this has been reported as open ticket attached.
The module was faulty we release it.
We noticed SSL BUG on the radios we have repaired this problem should now be working again.
The Incident was reported that DNS wasn't responding to the following requests.
Planned: We are going to add more DNS connection in-coming weeks.
Review: SDNS: connections: - We still reviewing the solutions in the best way forward.
Report : CDN - connection - bad match on SSL/TLS = miss match.
This was reported : was addressed in the patch update.
The customers website functions without any further issues.
We have reviewed all information about number of reports been sent or notifications. Please note we have reviewed all feedback. We plan to remove the plesk notification due to the complaints received.
Fixed : Notification : has been switched off due to the customer complaints about amount of notification received
The api sending fault or default page when customer using the services login api we're checking.
Fix : We are looking for fix for the Customer-Manager : We will fix these errors very soon.
Fixed API : is now connecting to the correct address : CLOSED : 17/01/2021
If you require any support contact customer care!
We are making some changes via our CDN - Networks Protocols today 15/01/2021
These changes include control panel might show errors on our CDN - Networks we are process changes via the bots.
Updates will be provided shortly
These are active we are checking further make sure working :
These connection now online : We are now checking the API connection towards customer account > to the back-end connections
Customer " Reported " control-panel was running slow on response
This was due to error in the API and redirect to the CDN
We noticed packet drops detected on interfaces. We going are going change Network connector.
Reason for closing : Date 2020-12-23 06:14:09 GMT (UTC +00:00), Network connector:
We have replaced the network cable.
We don't see any packet drops detected.
We are aware off an issue with the DNS Template this is whats causing customers websites not to load,
we are gonna be implementing a fix and will update once this has been implemented.
UPDATE: implemented a fix
REASON: implemented a fix and have closed this network issue down
We are aware of an issue with the client area loading slowly we are looking into this as we speak
We will update when we have implemented the fix
UPDATE:- implemented fix
Reason:- implemented fix has been done reason for closing.
Back-end login slow : Causing issues on the N-User login via API.
Update as of 11/12/2020 @ 18:29 we have implemented a fix on the back-end login and the N-user will notice that the login is a lot quicker
we are having with issues with automatic email config server settings failure we are fully aware of this and are investigating.
We apologise for this and will update yourselves once we have rectified this.
UPDATE as of 10/12/2020 @12.28 :- we have find the issue and it is related to pop3s we are gonna be implementing a fix as we speak and will update again shortly.
We plan Anti-Spam - for the customer's control own anti-spam filers system via the API - back-end functions.
This due to increase of large amounts of Spam.
|#JCQ-50098 - spam emails|
Intervention will be carried out on the 30th Nov 2020 at 12 am GMT - UTC +1 We noticed malfunction issues that have affected the entire of the cloud-InfraStructure.
Customer Solutions Teams & Intervention Teams.
NTP - Services now fixed.
- ScoobyNetworks is experiencing delays in updating customer SSL
We will update the status once we implement the fix.
- The issue has been identified and a fix is being implemented.
- This incident has been resolved.
We're making improvements to the Stunnel's on ScoobyNetworksGroup!
Intervention will be placed to fix some bugs
ScoobyNetworksGroup - Support
We're announcing we are going for full infrastructure reboot at 12'oclock tonight GMT on 20th June.
These are upgrading our systems and security patches.
this has been completed as of 2006/2020 @ 02.42
Just to make customers aware there maybe some downtime on the system as we doing updates
we will update shortly when these have been done
Update has now been completed on the 14/06/2020 at 19:57
We made further improvements to the firewalls at 02:00 O'clock GMT-UTC Time.
No further action is needed.
Closed : 13/06/2020 - 05:06am
We are aware of on-going issue customers unable to login in the manager.
Google Accounts - We are checking....
The following incident was rectify with the following fix at 07:42AM - On 11/06/2020. The API / htaccess was outdated with our following updates.
We are sorry for the inconvenience during the early hours of the Central United Kingdom (UK) time GMT-UTC.
If you experience any other following issues please don't hesitate to contact our customer care.
The incident has now been resolved at 07:33AM.
We are making progress on our latest dns-cluster's we are changing our dns communications via our networks online.
Our next step is to change our DNS services from online to co.uk
We have done some changes on our " Web-Hosting" control panel we updated from 7.3 to 7.4 since done this none of the website that use PHP will no longer display on Firefox or Google Chrome. We have implemented a fix and we are monitoring the situation.
DNS servers responded ERROR: One or more of your nameservers did not respond:
The ones that did not respond are:
We are going make some changes to the Customers - Control panels allow them to add RDNS in the control panels.
Task : Change completed at 20/05/2020
to deploy duo update to the entire network
To deploy sh script to auto install 3cx
A problem occurred and 3CX Phone System was not updated successfully. Your installation at
encountered a problem and the update was not successful. The system was rolled back to the previous version. Attached is the update log files which contain information about why the update failed.
UPDATE: the issue with the phone system has now be resolved.
Investigation : We are working to understand the full impact and mitigate this problem. More updates to follow shortly
Process : Investigating issues with network performance.
Monitoring - A fix has been implemented and we are monitoring the results.
We have troubles with our "message": "The requested object (ipMitigationProfile : does not exist :
We had API fault with processing orders.
Admin : This module and API now responding again.
We have added module update and edit the API function shouldn't have anymore troubles.
Our Administrator adding new features coming soon to our latest update of Manager.
We are aware of Module command error we have disabled this function.
Provide updates shortly.
This Module is fixed.
This intervention done.
effected network is the entire network ,we are currently mitigating large amount of data
mite experience jitter/lagg that we are in the process of switching the old gateways offline
we will update this status once mitigation and intervention start
Glenn system Administrator engineering team
We planned Maintenance at 12 GMT - 12-30 GMT the entire networks will be-offline.
We are changing our current status to on-hold until we have update from our engineers.
We're adding new firewall rules to our latest ssl-tunnel connections.
Update : The SSL-Tunnel will be removed from "Windows" moved to Cloud-Linux Services in process.
Update: We had some troubles with firewall rules we detection of attack on the s-tunnel (IP) address. This didn't cause any troubles with the radio streaming on the sites or app or other services.
Update: We now monitor for 24 hour's see if any problems. - We are not closing the task at the moment we will monitor until we're happy to close the task.
connection unable to connect we're working on fix.
Report has been made for faulty "MYSQL" services we will comment when we know more.
MYSQL - Services going for restart.
Repair has been made we're closing the following intervention.
We're will start intervention in about 15 minutes time to re-config Recommended values .
Thanks, Support Teams.
UPDATE : Description: A link between San Jose (SJC) et Singpore (SNG) is down, we are checking with provider for a quick resolution
Impact: Latency may increase by 5ms maximum for customers in north america
DNS Failed respond.
We are monitoring the systems. We have restarted the services.
We will bring WHMSonic support with SSL coming soon in 2020
Thank You - Dev Teams
Replicated issue on our side with connection to the "Central - Servers. Case #FWN-9821 is open with our developers in order to have this reviewed for future releases.
Administrator : ASA# sh int | i error 0 input errors, 0 CRC, 0 frame, 0 overrun, 0 ignored, 0 abort 0 output errors, 0 collisions, 0 interface resets.
05440947 packets input, 2283932667 bytes, 0 no buffer Received 1161 broadcasts, 0 runts, 0 giants, 0 throttles 0 input errors, 0 CRC, 0 frame, 0 overrun, 0 ignored 0 watchdog, 0 multicast, 0 pause input 0 input packets with dribble condition detected 132064543 packets output, 671431731 bytes, 0 underruns 27627 output errors, 22353 collisions, 3 interface resets 0 babbles, 27627 late collision, 66314 deferred 0 lost carrier, 0 no carrier, 0 PAUSE output 0 output buffer failures, 0 output buffers swapped out
In a normal TCP session, the client sends a SYN packet to the server, with the MSS included within the TCP options of the SYN packet; the server, upon receipt of the SYN packet, should recognize the MSS value sent by the client and then send its own MSS value in the SYN-ACK packet—once both the client and the server are aware of each other's MSS, neither peer should send a packet to the other that is greater than that peer's MSS Verify from syslogs that the flow is not failing due to MSS issues %ASA-4-419001: Dropping TCP packet from outside:172.18.25.60/443 to inside:10.110.61.153/61312, reason: MSS exceeded, MSS 1380, data 1460
FWSM-1# sh xlate debug | i 10.130.6.20 NAT from INSIDE:10.130.6.20 to OUTSIDE:10.130.6.20 flags Ii idle 0:00:42 timeout 3:00:00 connections 4 NAT from INSIDE:10.130.6.20 to WAN:10.130.6.20 flags Ii idle 1:55:26 timeout 3:00:00 connections 0 FWSM-1# sh xlate detail | i 10.101.2.128 NAT from TEST:10.101.2.128 to INSIDE:10.101.2.128 flags Ii
Firewall#sh local | i host|count/limit local host: <10.10.57.5>, TCP flow count/limit = 16/unlimited TCP embryonic count to host = 93404 UDP flow count/limit = 6/unlimited local host: <10.10.57.251>, TCP flow count/limit = 53/unlimited TCP embryonic count to host = 140948 UDP flow count/limit = 1/unlimited
Identifying Translation Problems—Traffic Breaks: ―Clear Local‖ Resolves the Problem Temporarily
Relevant config: interface Vlan102 nameif WAN security-level 10 ip address 10.130.4.139 255.255.255.248 interface Vlan319 nameif INSIDE security-level 80 ip address 10.130.1.129 255.255.255.248 no nat-control route INSIDE10.130.6.0 255.255.255.0 10.130.1.131 1 route WAN 10.101.0.0 255.255.0.0 10.130.4.141 1 host (10.130.6.20)—Router—(80)INSIDE-Vlan319/— FWSM—/Vlan102-WAN(10)—(WAN router)— (10.101.2.128)
failover lan unit primary failover lan interface FAILOVER GigabitEthernet0/3.2624 failover polltime unit msec 200 holdtime msec 800 failover polltime interface msec 500 holdtime 5 failover replication http failover link STATE GigabitEthernet0/2.2623 failover interface ip FAILOVER 10.10.11.37 255.255.255.252 standby 10.10.11.38 failover interface ip STATE 10.10.11.41 255.255.255.252 standby 10.10.11.42
The newly (failover) enabled primary unit will automatically detect the existing sec/act unit and will sync up with it ASA(config)# failover.
We have found the cause of the root of the problem we will investigate more.
We had report of Control Panel due to rules.
In order to have better routing for our internal traffic, we are going to change some configuration of our InterDC infrastructure.
No impact is expected.
Customer "REPORTED" on-going issues with "RADIO" stream we' adding some new "RULES" filter out the connection.
Mod Added we're monitoring connections.
: checked done. This task still open so we might make new adjustments.
We plan to bring "Special" services online we' believe that (BACKUPS) very important so we are working together with cPanel to provide us new services.
ScoobyNetworks Customer Services Support!
We're are aware of problem affecting WHMSonic access control panels we're working around the clock to resolve issue.
The issue is out of our control we're hope have update for you on 28th Nov 2019. If you have any issues please contact support.
ScoobyNetworks Cpanel Teams!
We're are aware of problem with controls - settings it seems like our API is at fault we're checking we're provide more updates shortly.
the control/settings are now working.
Updated version will be installed in about 20 minutes time intervention is set.
Our incident teams are changing the network card in this server.
intervention is set start in 20 minutes.
ScoobyNetworks. Incident Teams.
#fixed changed done.
We're in process of rolling out final build of Firewall Rules on our Network.
after long months we've finally cracked our rules and able to provide better secure services.
Please stand-by for updates!
Updates : We' are changing some packet filters today!
# Updates : We're starting next process of the firewall in-coming weeks
#Postponed we're provide more updates at later date.
We are working on our LDN<>AMS darkfiber.
The goal is to add a RAMAN COP on the submarine span.
Update : #Services/Network - Process!
Impact / Affected perimeter: instances
• Impact type / Impact type: downtime
• Estimated time to recovery / Estimated resolution time: NO ETA minutes
We' had reports some customers unable to access systems at the present moment we're checking we will provide updates shortly!#
We have latency problems affecting entire network, we will update it as soon as we get response from the network team.
We have detected an increase of latency on all the traffic passing trough by _eri1 WESTERN Europe.
We occurred some instabilities, we attempted re-routing process traffic will flow CENTRAL Europe, We sincerely contact our upstream providers for further questions.
We're just updating Software we're be back soon.
ScoobyNetworks Group Support!
We are having and issue with hardware please bare with us as we are getting this sorted in next 24 to 48 hours.
This will include clients control panels and radio control panels. This also includes now password reset been fixed on login area.
Please bare with us as these issues will be sorted in 24 to 48 hours ( SERVICE COULD BE INTERRUPTED. )
Due to these issues we are aware of the issues we are having. if you would like to contact us please call us on out numbers.
THANKS BILLING DEPARTMENTS
update : fixed
We're noticed miss-configuration has been done on some rules we're working on these re-correct errors.
you may having issues connecting to services whiles this work is carried out.
Update : We're noticed some of the rules aren't right we're checking.
Update : Rules are now in place we're happy to confirm end of intervention.
Thank you ScoobyNetworks Anti Ddos Teams.
adjusted firewall rules and filters.
we're should be done in the next 1 hours time please standby for updates.
Closed : 1:59 PM done!
Intervention : We're checking we're unable to connect to the bot network.
Intervention : We have rebooted server everything back online.
We're having some troubles connecting via the backup managers.
backup-manager hasn't reported any data since 7/7/19 we're investigating
- We're found the root cause of the backup manager not connecting we're implementing fix at 7:48 AM.
We're provide more updates shortly.
implementing the fix server ping ok and on login screen.
Reason for closing : Done
REPORTS : 100% CPU usage.
we're checking we're provide updates shortly.
Stable, no further input.
2019 order process will start here intervention will happen shortly.
stand-by for agent to reply on the on-going process.
you will not find any IP address or any data here this just records of the process when the process was done and how long when it was closed.
Start : Windows Process *STARTED 9:35 US E-Time
Install : *Hidden*
Refresh : Hidden*
Login : Hidden*
End : Hidden*
2019 server process was install ping online working okay done.
Dropping connection - Investigating.
08/06/2019 - change was made to firewall - monitor connection
09/06/2019 - Firewall re-config - monitor connection for further 12 hours.
09/06/2019 - module was affected - reinstalled module connection stable.
Information : reports of GL-SNS was affected in connection drop from the last 24 hrs we're investigated re-installed fault module. /
We're check connection for next 12 hrs.
10/06/2019 - 2:40 AM - change of IP address - we're still checking see if connection stable.
Report : opened internal report with provider.
on-going - no reply back from provider or answer from their network.
on 11/06/2019 - reply, provider we're corrected the problem all should be stable.
on 11/06/2019 - we closed case.
more information - https://status.scoobynetworks.uk/782558560
This isn't greats news's that our teams going to announce. Our Administrators have been working tireless on Cloud despaired off the network chart this morning. Alert was sent to Administrators.
We looked at the remote wasn’t any access. We connected to it with our IPMI console. We found the system in emergency mode. We have been recovering all data since this morning our teams have been none stop. We’re still investigating the course of the root problem. We’re sorry we can’t give any information at this time.
Case ID 16/05/2019-Authentication.
Administrator : interv - gain access via SSH.
* Authentication error - stopped N-users unable to access Services/Servers.
Reason : This was due NTP SERVER having the in-correct time.
Closed:16/05/2019 - 4:48 PM
We're aware of issue causing the bot's from disconnect we're working on fix
We're monitor connection for 1 hour provide updates shortly
Update : 03:20 AM 12/05/2019 - We've found the cause of the problem causing connection to disconnect we're stable the connections we're online i'm closing this case.
We're investigating dropping connection towards the bot central networks we're provide updates shortly.
Update : 2:36 AM - Administrator - We understanding some Customer/Client's notice current bots disconnecting our Administrator currently still checking and monitoring the connections.
please don't use any commands to pause or stop the bot's at this time. Please don't make any-changes via your bot at this moment your changes might not take affect. Please wait for the next update within 24 hours time.
Our Administrator, re-called the CASE now resolved any further information will not be provided as this CASE now resolved
named servers and records added SRV - that caused 3 second delay in the network we're now fixed this problem.
SRV - Records - Added fixed.
Closed no further action.
We're making some changes to our network and updating security and some patches.
ScoobyNetworks Group Teams.
ScoobyNetworks Group, restart successful - resolved.